# Img-Diff: Contrastive Data Synthesis for Multimodal Large Language Models

### v1

@@ Abstract.
High-performance Multimodal Large Language Models (MLLMs) rely heavily on data quality. This study introduces a novel dataset named Img-Diff, designed to enhance fine-grained image recognition in MLLMs by leveraging insights from contrastive learning and image difference captioning. By analyzing object differences between similar images, we challenge models to identify both matching and distinct components. We utilize the Stable-Diffusion-XL model and advanced image editing techniques to create pairs of similar images that highlight object replacements. Our methodology includes a Difference Area Generator for object differences identifying, followed by a Difference Captions Generator for detailed difference descriptions. The result is a relatively small but high-quality dataset of “object replacement” samples. We use the the proposed dataset to finetune state-of-the-art (SOTA) MLLMs such as MGM-7B, yielding comprehensive improvements of performance scores over SOTA models that trained with larger-scale datasets, in numerous image difference and Visual Question Answering tasks. For instance, our trained models notably surpass the SOTA models GPT-4V and Gemini on the MMVP benchmark. Besides, we investigate alternative methods for generating image difference data through “object removal” and conduct thorough evaluation to confirm the dataset’s diversity, quality, and robustness, presenting several insights on synthesis of such contrastive dataset. To encourage further research and advance the field of multimodal data synthesis and enhancement of MLLMs’ fundamental capabilities for image understanding, we release our codes and dataset at https://github.com/modelscope/data-juicer/tree/ImgDiff.
@@
Figure 1. Examples of generated “object replacement” data within Img-Diff.
@@ 1. Introduction
The emergence of large language models (LLMs) has revolutionized natural language processing (NLP) (Mann et al., 2020; Touvron et al., 2023). This advancement has paved the way for the development of Multimodal Large Language Models (MLLMs) that seamlessly integrate linguistic and visual understanding. Improving the performance of MLLMs hinges on two primary avenues: evolving model architectures and enhancing dataset quality. The majority of state-of-the-art MLLMs (Liu et al., 2023c, b, 2024; Bai et al., 2023; Li et al., 2023d) implement a two-phase approach, commencing with a pretraining phase involving extensive image-text pairs for modality alignment, followed by fine-tuning aimed at optimizing visual question answering (VQA) capabilities with specific instruction tuning datasets.
The efficacy of pretraining datasets profoundly affects MLLMs’ capabilities in performing core visual tasks. Concurrently, the quality of visual instruction tuning datasets plays a crucial role in the overall performance of MLLMs in VQA tasks and diverse downstream applications. With the evolution of visual instruction tuning datasets, several recent studies have successfully integrated object detection and Optical Character Recognition (OCR) datasets, such as Refcoco (Kazemzadeh et al., 2014), Visual Genome (Krishna et al., 2017), OCR-VQA (Mishra et al., 2019), and TextVQA (Singh et al., 2019), significantly enhancing MLLMs’ proficiency in tasks requiring detailed image perception.
In light of advancements in contrastive learning and image difference captioning research in specific domains (Black et al., 2024; Park et al., 2019; Jhamtani and Berg-Kirkpatrick, 2018), which underscores the potential of pair of different images in refining models’ image recognition capabilities, we introduce a general-purpose yet challenging dataset named Img-Diff. This novel dataset sets itself apart from existing visual instruction tuning datasets by generating pairs of very similar images featuring subtle object alterations. Rather than compelling MLLMs to focus solely on a single image, our dataset challenges them to analyze paired images and articulate the differences within designated regions. By doing so, we aim to empower MLLMs with enhanced abilities to scrutinize intricate areas within images and assess discrepancies across corresponding regions, thereby augmenting their fine-grained image recognition capabilities.
@@
We further evaluate the diversity and quality of our dataset, ensuring it encompasses a broad array of object categories while showcasing rich variability. Through meticulous manual labeling, we affirm the high quality of the dataset. Additionally, we conduct ablation studies to examine the effects of various filtering intensities,. We also investigate an alternative methodology for constructing image difference data focused on “object removal”, assessing its effectiveness on LLaVA-1.5-7B and MGM-7B, and present fruitful insights on construction of contrastive data synthesis.
@@
  • We present a novel data synthesis method and an effect-proven Img-Diff dataset, comprising pairs of very similar images generated through object replacement, with a focus on processes such as segmentation, filtering, and detailed captioning of image difference .
  • We execute visual instruction tuning on LLaVA-1.5-7B and MGM-7B using our dataset, and rigorously assess the fine-tuned models’ performance on many image difference benchmarks and widely-used MLLM benchmarks, demonstrating substantial performance improvements.
  • We provide a comprehensive evaluation of the diversity and quality of the proposed dataset, confirming its richness and high standard. Through ablation studies, we identify empirical good filtering thresholds for such contrastive dataset.
  • We open-source our dataset and codes at () to facilitate ongoing research, encouraging innovative endeavors in more MLLM datasets and image difference methodologies.
@@ 2.1. Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements since their introduction. Research highlights two key factors that primarily influence the effectiveness of MLLMs: model architecture and dataset quality (Qin et al., 2024).
With respect to model architecture, notable approaches include Flamingo (Alayrac et al., 2022), IDEFICS (Hugo et al., 2023; Laurençon et al., 2024), BLIP-2 (Li et al., 2023b), and Qwen-VL (Bai et al., 2023), which leverage learnable queries to extract essential information from the CLIP (Radford et al., 2021; Dosovitskiy et al., 2020) image features. Alternatively, LLaVA (Liu et al., 2023c, b, 2024) and MGM (Li et al., 2023d) utilize projection-based interfaces to facilitate interactions between text and image modalities. Furthermore, LLaMA-Adapter (Zhang et al., 2023) and LaVIN (Luo et al., 2023) implement parameter-efficient tuning mechanisms to transfer image-related information to the LLM. A recent work also show usefullness of object detection model for MLLMs (Jiao et al., 2024).
From the perspective of datasets, there are two prevalent strategies: enhancing the quality of pretraining data and improving visual instruction tuning data. The former aims for better semantic alignment between images and text by introducing substantial numbers of image-text pairs, enabling MLLMs to proficiently address fundamental visual tasks, such as image captioning. However, MLLMs trained exclusively on pretraining datasets may encounter difficulties with VQA challenges. Recent research has increasingly concentrated on refining visual instruction tuning datasets, enabling MLLMs to enhance performance across diverse question-answering tasks. Works like LLaVA (Liu et al., 2023c, b, 2024), InstructBLIP (Dai et al., 2023), SPHINX (Lin et al., 2023), and MGM (Li et al., 2023d) leverage high-quality finetuning datasets characterized by extensive task diversity, allowing models to excel in tasks related to image perception, reasoning, and optical character recognition (OCR). Additionally, methods such as Shikra (Chen et al., 2023), ASM (Wang et al., 2023), and PINK (Xuan et al., 2023) utilize substantial amounts of object detection data to enhance the models’ localization capabilities.
@@ 2.2. Datasets on Image Differences
Datasets focused on image differences typically consist of pairs of similar images supplemented with textual descriptions of their variations. For instance, the Spot-the-Diff dataset (Jhamtani and Berg-Kirkpatrick, 2018) contains pairs of street scenes captured at different times by the same surveillance cameras. The CLEVR-Change dataset (Park et al., 2019) delineates scene variations of geometric objects against a clean backdrop. The Birds-to-Words dataset (Wah et al., 2011) elaborates on the nuanced differences among various bird species found in natural habitats. The Image-Edit-Request (Tan et al., 2019) dataset features edited images alongside their original counterparts, accompanied by descriptions of the modifications made.
@@ 3.1. Overview
In recent years, contrastive learning methods have significantly improved the image-text understanding of vision-language models(Radford et al., 2021). These methods involve repeatedly pairing batches of images and texts, requiring the model to distinguish between matching and non-matching image-text pairs, which enhances the model’s capability to differentiate between semantically similar and dissimilar image-text pairs. Our method incorporates contrastive learning principles and applies them to generate MLLM image-text data. Specifically, our method focuses on replacing objects between the image pairs, requiring MLLMs to identify similarities and differences in specific areas of the images, which aims to improve MLLMs’ abilities to recognize fine-grained differences in images.
The process of generating “object replacement” data can be divided into three parts. The first part is to generate similar images and form image pairs, where the only difference between these pairs are object replacement. Next, we name the second part “Difference Area Generator”, which extracts bounding box regions containing object differences between the image pairs. The third part, called the “Difference Captions Generator”, uses an MLLM to generate descriptive text for the areas containing object differences and then create question-answer pairs with questions such as “What objects have changed in this area?”. The process involves many model selections, which are discussed in Appendix B.2. The process also involves multiple filtering operations, with the specific thresholds detailed in Appendix B.3, which we determine through experimental comparisons in Section 4.6. The sample examples are presented in Appendix C. The general framework is shown in Figure 2. To enhance the reproducibility and reusability, our proposed components and the end-to-end construction pipeline are implemented as data processing operators and configurable data recipes of Data-Juicer (Chen et al., 2024a, b) respectively.
@@ 3.2. Image Pairs Generation
Before generating the Img-Diff data, we first need to generate numerous pairs of similar images. The process is shown in Figure 2, which involves replacing the objects in images. We refer to InstructPix2Pix(Brooks et al., 2023) and employ a generative model combined with image editing techniques to generate image pairs. However, while InstructPix2Pix utilizes the Stable-Diffusion-1.5, we use the more advanced Stable-Diffusion-XL(Podell et al., 2023) to produce more realistic images.
We start by obtaining 118K image captions from MS COCO(Lin et al., 2014), which are descriptions biased towards real photos. Then, we use the LLM Vicuna-1.5-13B(Chiang et al., 2023) to perform object replacement in the captions. The prompt used is “Here is a sentence: ‘INPUT’. Please only replace one of the objects in this sentence with another object.” Here, INPUT refers to the original caption, and the answers from the LLM are the new captions. Finally, based on the caption pairs, we use the text-to-image generative model Stable-Diffusion-XL and image editing technique Prompt-to-Prompt(Hertz et al., 2022) to generate image pairs with only few objects replaced.
@@ 3.3.1. Overview
The Difference Area Generator aims to identify the locations of object differences between the image pair. Although object detection models are capable of identifying objects inside images, the range of object categories is quite limited(Zhang et al., 2022; Jocher et al., 2023). Therefore, to increase the number of detectable object categories and enhance dataset diversity, we develop the Difference Area Generator based on segmentation and image similarity comparisons. The process is illustrated in Figure 3.
In this phase, we first use the Image Similarity Filter to obtain image pairs with high similarity but not completely identical and use FastSAM(Zhao et al., 2023) to perform image segmentation on each image. Next, we crop the images based on the bounding box information got from segmentation and use the Image-text Matching Filter to filter the cropped sub-images for the presence of valid objects. Finally, we use the Difference Detector to determine whether there are indeed differences between the bounding box regions of the image pairs and perform IoU (Intersection over Union) filtering to remove overlapping bounding boxes, ultimately obtaining valid bounding box information.
@@ 3.3.4. Difference Detector
The Difference Detector is used in the final stage of the Difference Area Generator. This module determines whether there are differences between the bounding box regions of the image pair. For a given bounding box, we first crop two sub-images from both image A and B based on it. These two sub-images are then filtered through the Image Similarity Filter and the bounding box is considered effective only if the difference is significant enough. After processing all bounding boxes, we use the IoU method to filter out overlapping bounding boxes. Only the bounding boxes with a higher degree of difference are retained. The remaining bounding boxes are the effective bounding boxes ultimately outputted by the Difference Area Generator.
@@ 3.4.1. Overview
After obtaining the valid bounding box regions, we use the Difference Captions Generator to generate difference captions about the content of these regions (with each round of process focusing on only one bounding box in one image pair). The reason our difference captions only target specific regions is that an image pair can contain multiple object differences and a single difference caption cannot fully capture all of them. Therefore, we highlight specific regions with red boxes and provide targeted difference captions to ensure greater accuracy.
@@ 3.4.2. Stage1: Object Labeling & Filter
In Stage1, for each image pair, we first select N bounding box regions with the lowest similarity between images (N is set to 5 in this project) as candidate regions. Then, for each bounding box, we use the MLLM LLaVA-NEXT(Liu et al., 2024) to describe its corresponding regions and then apply two filtering processes: the first filter is the Image-text Matching Filter, which checks whether the content of the regions matches the captions; the second filter is the Captions Similarity Filter, which assesses whether there are differences between the two captions. Once the filtering is complete, we obtain valid bounding boxes and captions for subsequent difference captioning.
@@ 3.4.4. Stage2: Difference Captions Generating
In Stage2, for each valid bounding box of each image pair, we first draw two red boxes into the images based on the bounding box information, highlighting the differences for easier localization. Then, we provide the MLLM LLaVA-NEXT with the captions of the bounding box regions and instruct it to describe the differences based on the content captions and the highlighted images. Finally, we can obtain the difference caption for the bounding box of the image pair.

### v2

@@ Abstract.
High-performance Multimodal Large Language Models (MLLMs) rely heavily on data quality. This study introduces a novel dataset named Img-Diff, designed to enhance fine-grained image recognition in MLLMs by leveraging insights from contrastive learning and image difference captioning. By analyzing object differences between similar images, we challenge models to identify both matching and distinct components. We utilize the Stable-Diffusion-XL model and advanced image editing techniques to create pairs of similar images that highlight object replacements. Our methodology includes a Difference Area Generator for object differences identifying, followed by a Difference Captions Generator for detailed difference descriptions. The result is a relatively small but high-quality dataset of “object replacement” samples. We use the the proposed dataset to finetune state-of-the-art (SOTA) MLLMs such as MGM-7B, yielding comprehensive improvements of performance scores over SOTA models that trained with larger-scale datasets, in numerous image difference and Visual Question Answering tasks. For instance, our trained models notably surpass the SOTA models GPT-4V and Gemini on the MMVP benchmark. Besides, we investigate alternative methods for generating image difference data through “object removal” and conduct a thorough evaluation to confirm the dataset’s diversity, quality, and robustness, presenting several insights on the synthesis of such a contrastive dataset. To encourage further research and advance the field of multimodal data synthesis and enhancement of MLLMs’ fundamental capabilities for image understanding, we release our codes and dataset at https://github.com/modelscope/data-juicer/tree/ImgDiff.
@@
Figure 1. Three “object replacement” examples within Img-Diff, highlighting fine-grined difference in both vision and language.
@@ 1. Introduction
The emergence of large language models (LLMs) has revolutionized natural language processing (Mann et al., 2020; Touvron et al., 2023), also paved the way for the development of Multimodal Large Language Models (MLLMs) that seamlessly integrate linguistic and visual understanding. Improving the performance of MLLMs hinges on two primary avenues: evolving model architectures and enhancing dataset quality (Qin et al., 2024). The majority of state-of-the-art (SOTA) MLLMs (Liu et al., 2023c, b, 2024b; Bai et al., 2023; Li et al., 2023d) employ a two-phase approach, commencing with a pretraining phase involving extensive image-text pairs for modality alignment, followed by fine-tuning aimed at optimizing visual question answering (VQA) capabilities with specific instruction tuning datasets.
The efficacy of pretraining datasets profoundly affects MLLMs’ capabilities in performing core visual tasks. Concurrently, the quality of visual instruction tuning datasets plays a crucial role in the overall performance of MLLMs in VQA tasks and diverse downstream applications. With the evolution of visual instruction tuning datasets, several recent studies have explored integration of object detection and Optical Character Recognition (OCR) datasets, such as Refcoco (Kazemzadeh et al., 2014), Visual Genome (Krishna et al., 2017), OCR-VQA (Mishra et al., 2019), and TextVQA (Singh et al., 2019), effectively enhancing MLLMs’ proficiency in tasks requiring detailed image perception.
In this paper, we focus on a new direction for enhancing MLLM datasets, driven by the potential of pairs of different images to refine models’ image recognition capabilities, as demonstrated by advancements in contrastive learning and image difference captioning for conventional vision models (Zhang et al., 2023b; Black et al., 2024; Park et al., 2019; Jhamtani and Berg-Kirkpatrick, 2018). Specifically, we introduce a general-purpose yet challenging dataset named Img-Diff, which sets itself apart from existing visual instruction tuning datasets of MLLMs by generating pairs of highly similar images featuring subtle object alterations. Rather than compelling MLLMs to focus solely on a single image, our dataset challenges them to analyze paired images and articulate the differences within designated regions, meanwhile taking the high-quality textual descriptions of the difference as learning signals. By doing so, we aim to empower MLLMs with enhanced capabilities of fine-grained image recognition, via scrutinizing intricate areas within images and assess discrepancies across corresponding regions.
@@
We further evaluate the diversity and quality of our dataset, ensuring it encompasses a broad array of object categories while showcasing rich variability. Through meticulous manual labeling, we affirm the high quality of the dataset. Additionally, we conduct ablation studies to examine the effects of various filtering intensities. We also investigate an alternative methodology for constructing image difference data focused on “object removal”, assessing its effectiveness on LLaVA-1.5-7B and MGM-7B, and present fruitful insights on the construction of contrastive data synthesis.
@@
  • We present a novel data synthesis method and an effect-proven Img-Diff dataset, comprising pairs of highly similar images, with a focus on processes such as segmentation, filtering, and detailed captioning of image difference.
  • We execute visual instruction tuning on LLaVA-1.5-7B and MGM-7B using our dataset, and rigorously assess the fine-tuned models’ performance on many widely-used MLLM benchmarks, demonstrating substantial performance improvements.
  • We provide a comprehensive evaluation of the diversity and quality of the proposed dataset, confirming its richness and high standard. Through ablation studies, we identify good empirical filtering thresholds for such contrastive datasets.
  • To facilitate ongoing research, encouraging innovative endeavors in more MLLM datasets and image difference methodologies, we open-source our dataset and codes at https://github.com/modelscope/ data-juicer/tree/ImgDiff.
@@ 2.1. Multimodal Large Language Models
MLLMs have exhibited remarkable advancements and continue to evolve rapidly (Achiam et al., 2023; Liu et al., 2024b, a; Google, 2023). Research highlights two key factors that primarily influence the effectiveness of MLLMs: model architecture and dataset quality (Qin et al., 2024).
With respect to model architecture, notable approaches include Flamingo (Alayrac et al., 2022), IDEFICS (Hugo et al., 2023; Laurençon et al., 2024), BLIP-2 (Li et al., 2023b), and Qwen-VL (Bai et al., 2023), which leverage learnable queries to extract essential information from the CLIP (Radford et al., 2021; Dosovitskiy et al., 2020) image features. Alternatively, LLaVA (Liu et al., 2023c, b, 2024b) and MGM (Li et al., 2023d) utilize projection-based interfaces to facilitate interactions between text and image modalities. Furthermore, LLaMA-Adapter (Zhang et al., 2023a) and LaVIN (Luo et al., 2023) implement parameter-efficient tuning mechanisms to transfer image-related information to the LLM. A recent work also shows the usefulness of the object detection model for MLLMs (Jiao et al., 2024).
From the perspective of datasets, there are two prevalent strategies: enhancing the quality of pretraining data and improving visual instruction tuning data. The former aims for better semantic alignment between images and text by introducing substantial numbers of image-text pairs, enabling MLLMs to proficiently address fundamental visual tasks, such as image captioning. However, MLLMs trained exclusively on pretraining datasets may encounter difficulties with VQA challenges. Recent research has increasingly concentrated on refining visual instruction tuning datasets, enabling MLLMs to enhance performance across diverse question-answering tasks. Works like LLaVA (Liu et al., 2023c, b, 2024b), InstructBLIP (Dai et al., 2023), SPHINX (Lin et al., 2023), and MGM (Li et al., 2023d) leverage high-quality finetuning datasets characterized by extensive task diversity, allowing models to excel in tasks related to image perception, reasoning, and optical character recognition (OCR). Additionally, methods such as Shikra (Chen et al., 2023), ASM (Wang et al., 2023), and PINK (Xuan et al., 2023) utilize substantial amounts of object detection data to enhance the models’ localization capabilities.
@@ 2.2. Datasets on Image Differences
Datasets focused on image differences typically consist of pairs of similar images supplemented with textual descriptions of their variations. For instance, the Spot-the-Diff dataset (Jhamtani and Berg-Kirkpatrick, 2018) contains pairs of street scenes captured at different times by the same surveillance cameras. The CLEVR-Change dataset (Park et al., 2019) delineates scene variations of geometric objects against a clean backdrop. The Birds-to-Words dataset (Wah et al., 2011) elaborates on the nuanced differences among various bird species found in natural habitats. The Image-Edit-Request (Tan et al., 2019) dataset features edited images alongside their original counterparts, accompanied by descriptions of modifications made.
@@ 3.1. Overview
In recent years, contrastive learning have significantly enhanced the image-text understanding of vision-language models (Radford et al., 2021; Zhang et al., 2023b). These methods typically involve constructing batches of images and texts, requiring the model to distinguish between matching and non-matching image-text pairs, which improves its ability to differentiate between semantically similar and dissimilar pairs.
Our approach leverages the principles of contrastive learning to generate MLLM image-text data. Specifically, it focuses on replacing objects within image pairs, encouraging MLLMs to identify similarities and differences in specific regions. This methodology aims to enhance the models’ capacity to recognize fine-grained differences in images, guided by textual descriptions that highlight detailed distinctions.
The process of generating “object replacement” data can be divided into three key parts. The first part involves creating similar images and forming image pairs, where the only difference between these pairs is the object replacement (Section 3.2). The second part, referred to as the “Difference Area Generator”, extracts bounding box regions that contain object differences between the image pairs (Section 3.3). The third part, termed as the “Difference Captions Generator” (Section 3.4), utilizes MLLMs to generate descriptive text for the areas with object differences and creates question-answer pairs, such as “What objects have changed in this area?”
This process incorporates multiple filtering operations, with specific thresholds outlined in Appendix B.4, which we determine through experimental comparisons in Section 4.6. Additional experimental details, including model selection and the training process, are also presented in Appendix B. For readability, we present several specific data examples in Appendix C and illustrate the general framework in Figure 2.
To enhance reproducibility and reusability, our proposed components and end-to-end construction workflow are implemented as data processing operators and configurable data recipes within Data-Juicer (Chen et al., 2024a, b).
@@ 3.2. Image Pairs Generation
Firstly, we need to generate numerous pairs of similar images as candidates. The process is shown in Figure 2, which involves replacing the objects in images. We refer to InstructPix2Pix (Brooks et al., 2023) and employ a generative model combined with image editing techniques to generate image pairs. However, while InstructPix2Pix utilizes the Stable-Diffusion-1.5, we use the more advanced Stable-Diffusion-XL (Podell et al., 2023) to produce more realistic images.
We start by obtaining 118K image captions from MS COCO (Lin et al., 2014), which are descriptions biased towards real photos. Then, we use the LLM Vicuna-1.5-13B(Chiang et al., 2023) to perform object replacement in the captions. The prompt used is “Here is a sentence: ‘INPUT’. Please only replace one of the objects in this sentence with another object.” Here, INPUT refers to the original caption, and the answers from the LLM are the new captions. Finally, based on the caption pairs, we use the text-to-image generative model Stable-Diffusion-XL and the image editing technique Prompt-to-Prompt(Hertz et al., 2022) to generate image pairs with only a few objects replaced.
@@ 3.3.1. Overview
The Difference Area Generator aims to identify the locations of object differences between the image pair. Although object detection models are capable of identifying objects inside images, the range of object categories is quite limited (Zhang et al., 2022; Jocher et al., 2023). Therefore, to increase the number of detectable object categories and enhance dataset diversity, we developed the Difference Area Generator based on segmentation and image similarity comparisons. The process is illustrated in Figure 3.
In this phase, we first use the Image Similarity Filter to obtain image pairs with high similarity but not completely identical and use FastSAM (Zhao et al., 2023) to perform image segmentation on each image. Next, we crop the images based on the bounding box information obtained from segmentation and use the Image-text Matching Filter to filter the cropped sub-images for the presence of valid objects. Finally, we use the Difference Detector to determine whether there are indeed differences between the bounding box regions of the image pairs and perform IoU (Intersection over Union) filtering to remove overlapping bounding boxes, ultimately obtaining valid bounding box information.
@@ 3.3.4. Difference Detector
The Difference Detector is used in the final stage of the Difference Area Generator. This module determines whether there are differences between the bounding box regions of the image pair. For a given bounding box, we first crop two sub-images from both images A and B based on it. These two sub-images are then filtered through the Image Similarity Filter and the bounding box is considered effective only if the difference is significant enough. After processing all bounding boxes, we use the IoU method to filter out overlapping bounding boxes. Only the bounding boxes with a higher degree of difference are retained. The remaining bounding boxes are the effective bounding boxes ultimately outputted by the Difference Area Generator.
@@ 3.4.1. Overview
After obtaining the valid bounding box regions, we use the Difference Captions Generator to generate difference captions about the content of these regions (with each round of the process focusing on only one bounding box in one image pair). The reason our difference captions only target specific regions is that an image pair can contain multiple object differences and a single difference caption cannot fully capture all of them. Therefore, we highlight specific regions with red boxes and provide targeted difference captions to ensure greater accuracy.
@@ 3.4.2. Stage1: Object Labeling & Filter
In Stage 1, for each image pair, we first select N bounding box regions with the lowest similarity between images (N is set to 5 in this project) as candidate regions. Then, for each bounding box, we use the MLLM LLaVA-NEXT (Liu et al., 2024b) to describe its corresponding regions and then apply two filtering processes: the first filter is the Image-text Matching Filter, which checks whether the content of the regions matches the captions; the second filter is the Captions Similarity Filter, which assesses whether there are differences between the two captions. Once the filtering is complete, we obtain valid bounding boxes and captions for subsequent difference captioning.
@@ 3.4.4. Stage2: Difference Captions Generating
In Stage 2, for each valid bounding box of each image pair, we first draw two red boxes into the images based on the bounding box information, highlighting the differences for easier localization. Then, we provide the MLLM LLaVA-NEXT with the captions of the bounding box regions and instruct it to describe the differences based on the content captions and the highlighted images. Finally, we can obtain the difference captions for the bounding box of the image pair.